# Multiple quantum exceptional, diabolical, and hybrid points in multimode bosonic systems: I. Inherited and genuine singularities

### v1

@@ V.1 Linear configuration
$\gamma_{+}$ $\alpha_{\mp}$ $\langle b_{2}\rangle$, $\langle b_{4}\rangle$ $\langle\bm{\hat{B}_{1}}\rangle$ 1 2x2 4x2 2x2 4x2
$\langle b_{3}\rangle$, $\langle b_{5}\rangle$ $\langle\bm{\hat{B}_{2}}\rangle$ 1
$-\alpha_{\mp}$ $\langle b_{2}^{\dagger}\rangle$, $\langle b_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{3}}\rangle$ 1 2x2 2x2
$\langle b_{3}^{\dagger}\rangle$, $\langle b_{5}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}}\rangle$ 1
0 $\langle b_{1}\rangle$ $\langle\bm{\hat{B}_{5}}\rangle$ 1 2x1 2x1 2x1 2x1
$\langle b_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{6}}\rangle$ 1
$2\gamma_{+}$ $2\alpha_{\mp}$ $\langle b_{2}^{2}\rangle$, $\langle b_{4}^{2}\rangle$ $\langle\bm{\hat{B}_{1}^{2}}\rangle$ 1 1x4 16x4 1x3 4x3
$\alpha_{-}+\alpha_{+}$ $\langle b_{2}b_{4}\rangle$ 2 +
$2\alpha_{\mp}$ $\langle b_{3}^{2}\rangle$, $\langle b_{5}^{2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x4 1x3
$\alpha_{-}+\alpha_{+}$ $\langle b_{3}b_{5}\rangle$ 2
$-2\alpha_{\mp}$ $\langle b_{2}^{\dagger 2}\rangle$, $\langle b_{4}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{3}^{2}}\rangle$ 1 1x4 1x3
$-\alpha_{-}-\alpha_{+}$ $\langle b_{2}^{\dagger}b_{4}^{\dagger}\rangle$ 2
$-2\alpha_{\mp}$ $\langle b_{3}^{\dagger 2}\rangle$, $\langle b_{5}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{4}^{2}}\rangle$ 1 1x4 1x3
$-\alpha_{-}-\alpha_{+}$ $\langle b_{3}^{\dagger}b_{5}^{\dagger}\rangle$ 2
$2\alpha_{\mp}$ $\langle b_{2}b_{3}\rangle$, $\langle b_{4}b_{5}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{2}}\rangle$ 2 2x4 1x4
$\alpha_{+}+\alpha_{-}$ $\langle b_{2}b_{5}\rangle$, $\langle b_{4}b_{3}\rangle$ 2
$-\alpha_{\mp}+\alpha_{\mp}$ $\langle b_{2}^{\dagger}b_{2}\rangle$, $\langle b_{4}^{\dagger}b_{4}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{1}}\rangle$ 2 2x4 1x4 6x4
$-\alpha_{\mp}+\alpha_{\pm}$ $\langle b_{2}^{\dagger}b_{4}\rangle$, $\langle b_{2}b_{4}^{\dagger}\rangle$ 2
$\alpha_{\mp}-\alpha_{\mp}$ $\langle b_{3}^{\dagger}b_{2}\rangle$, $\langle b_{5}^{\dagger}b_{4}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{1}}\rangle$ 2 2x4 1x4
$\alpha_{\pm}-\alpha_{\mp}$ $\langle b_{5}^{\dagger}b_{2}\rangle$, $\langle b_{3}^{\dagger}b_{4}\rangle$ 2
$\alpha_{\mp}-\alpha_{\mp}$ $\langle b_{2}^{\dagger}b_{3}\rangle$, $\langle b_{4}^{\dagger}b_{5}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{2}}\rangle$ 2 2x4 1x4
$\alpha_{\pm}-\alpha_{\mp}$ $\langle b_{2}^{\dagger}b_{5}\rangle$, $\langle b_{4}^{\dagger}b_{3}\rangle$ 2
$-\alpha_{\mp}+\alpha_{\mp}$ $\langle b_{3}^{\dagger}b_{3}\rangle$, $\langle b_{5}^{\dagger}b_{5}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{2}}\rangle$ 2 2x4 1x4
$-\alpha_{\mp}+\alpha_{\pm}$ $\langle b_{3}^{\dagger}b_{5}\rangle$, $\langle b_{5}^{\dagger}b_{3}\rangle$ 2
$-2\alpha_{\mp}$ $\langle b_{2}^{\dagger}b_{3}^{\dagger}\rangle$, $\langle b_{4}^{\dagger}b_{5}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{4}}\rangle$ 2 2x4 1x4
$-\alpha_{+}-\alpha_{-}$ $\langle b_{2}^{\dagger}b_{5}^{\dagger}\rangle$, $\langle b_{4}^{\dagger}b_{3}^{\dagger}\rangle$ 2
$\alpha_{\mp}$ $\langle b_{2}b_{1}\rangle$, $\langle b_{4}b_{1}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{5}}\rangle$ 2 8x2 16x2 4x2 8x2
$\langle b_{2}b_{1}^{\dagger}\rangle$ $\langle b_{4}b_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{6}}\rangle$ 2
$\langle b_{3}b_{1}\rangle$, $\langle b_{5}b_{1}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{5}}\rangle$ 2
$\langle b_{3}b_{1}^{\dagger}\rangle$ $\langle b_{5}b_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{6}}\rangle$ 2
$-\alpha_{\mp}$ $\langle b_{2}^{\dagger}b_{4}\rangle$, $\langle b_{4}^{\dagger}b_{1}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{5}}\rangle$ 2 8x2 4x2
$\langle b_{2}^{\dagger}b_{1}^{\dagger}\rangle$ $\langle b_{4}^{\dagger}b_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{6}}\rangle$ 2
$\langle b_{3}^{\dagger}b_{1}\rangle$, $\langle b_{5}^{\dagger}b_{1}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{5}}\rangle$ 2
$\langle b_{3}^{\dagger}b_{1}^{\dagger}\rangle$ $\langle b_{5}^{\dagger}b_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{6}}\rangle$ 2
0 $\langle b_{1}^{2}\rangle$, $\langle b_{1}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{5}^{2}}\rangle$, $\langle\bm{\hat{B}_{6}^{2}}\rangle$ 1 4x1 4x1 3x1 3x1
$\langle b_{1}^{\dagger}b_{1}\rangle$ $\langle\bm{\hat{B}_{6}\hat{B}_{5}}\rangle$ 2
@@ 0 $\langle b_{1}\rangle$ $\langle\bm{\hat{B}_{5}}\rangle$ 1 2x1 2x1 2x1 2x1
Table 5: Real and imaginary parts of the complex eigenfrequencies $\bm{\Lambda_{j}^{\rm r}}-i\bm{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(5)}_{\rm l}}$, given in Eq. (69), for the five-mode linear bosonic system derived from the equations for the FOMs up to second order. For details, see the caption to Tab. 1.
V.2 Pyramid configuration
$\hat{H}_{5,{\rm p}}$ $=$ $\left[\hbar\epsilon\hat{a}_{1}^{\dagger}\hat{a}_{2}+\hbar\epsilon\hat{a}_{1}^{\dagger}\hat{a}_{4}+\hbar\epsilon\hat{a}_{1}^{\dagger}\hat{a}_{5}+\hbar\epsilon\hat{a}_{2}^{\dagger}\hat{a}_{3}+\hbar\epsilon\hat{a}_{2}^{\dagger}\hat{a}_{5}\right.$ (75)
$+\hbar\epsilon\hat{a}_{3}^{\dagger}\hat{a}_{4}+\hbar\epsilon\hat{a}_{3}^{\dagger}\hat{a}_{5}+\hbar\epsilon\hat{a}_{4}^{\dagger}\hat{a}_{5}+\hbar\kappa\hat{a}_{1}\hat{a}_{2}$
$+\hbar\kappa\hat{a}_{1}\hat{a}_{4}+\hbar\kappa\hat{a}_{1}\hat{a}_{5}+\hbar\kappa\hat{a}_{2}\hat{a}_{3}+\hbar\kappa\hat{a}_{2}\hat{a}_{5}$
[…]
The Hamiltonian $\hat{H}_{5,{\rm p}}$ given in Eq. (75) leads to the Heisenberg-Langevin equations (63) in which the dynamical matrix $\bm{M^{(5)}_{\rm p}}$ attains the form using the $2\times 2$ submatrices introduced in Eq. (10):
Motivated by the $\mathcal{PT}$ symmetry we assume:
$2\gamma_{5}=\gamma_{12}+\gamma_{34}.$ (83)
Under these conditions, diagonalization of the matrix $\bm{M^{(5)}_{\rm p}}$ leaves us the following five eigenvalues ${\lambda_{j}^{M^{(5)}_{\rm p}}}$:
$\lambda_{1}^{M^{(5)_{\rm p}}}$ $=$ $-i\gamma_{+},$
$\lambda_{2,3}^{M^{(5)}_{\rm p}}$ $=$ $-i\gamma_{+}-\xi\pm\beta_{1},$
$\lambda_{4,5}^{M^{(5)}_{\rm p}}$ $=$ $-i\gamma_{+}+\xi\pm\beta_{2}.$ (84)
The corresponding eigenvectors are obtained as follows:
$\bm{y_{4}^{M^{(5)}_{\rm p}}}$ $=$ $\left[\sigma_{+},\sigma_{+},\sigma_{+}^{8},\sigma_{+}^{*},1\right]^{T},$
Contrary to the eigenvalues of the models discussed above, the eigenvalues $\lambda_{j}^{M^{(5)}_{\rm p}}$ for $j=2,\ldots,5$ in Eq. (85) exhibit the linear dependence on $\xi$. This means that the DD inherited to all the above-discussed models is removed in this model. It remains only for the eigenvalue $\lambda_{1}^{M^{(5)}_{\rm p}}$ that, however, has no ability to form EDs.
$\frac{\kappa^{2}}{\epsilon^{2}}+\frac{\gamma_{-}^{2}}{5\epsilon^{2}}=1.$ (86)
The real parts of the eigenvalues $\lambda_{j}^{M^{(5)}_{\rm p}}$ forming QEPs ($j=2,\ldots,5$) are drawn in Fig. 2(g,h).
The eigenvalues and eigenvectors of the $10\times 10$ matrix $\bm{M^{(5)}_{\rm p}}$ formed from the expressions given in Eqs. (84) and (85) and also in Eqs. (13) and (14) (for details, see Appendix A) allow us to analyze the system dynamics via a diagonal dynamical matrix. The appropriate operators $\hat{\bm{b}}=\left[\hat{b}_{1},\hat{b}_{1}^{\dagger},\hat{b}_{2},\hat{b}_{3},\hat{b}_{3}^{\dagger},\hat{b}_{2}^{\dagger},\hat{b}_{4},\hat{b}_{5},\hat{b}_{5}^{\dagger},\hat{b}_{4}^{\dagger}\right]^{T}$ then allow to identify the QEPs and QHPs observed in the dynamics of the first- and second-order FOMs, as summarized in Tab. 6).

### v3

@@ 0 $\langle\hat{b}_{1}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x9 1x6
$\gamma_{+}$ $\alpha_{\mp}$ $\langle\hat{b}_{2}\rangle$, $\langle\hat{b}_{4}\rangle$ $\langle\bm{\hat{B}_{1}}\rangle$ 1 2x2 4x2 2x2 4x2
$\langle\hat{b}_{3}\rangle$, $\langle\hat{b}_{5}\rangle$ $\langle\bm{\hat{B}_{2}}\rangle$ 1
$\alpha_{\mp}$ $\langle\hat{b}_{2}^{\dagger}\rangle$, $\langle\hat{b}_{4}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{3}}\rangle$ 1 2x2 2x2
$\langle\hat{b}_{3}^{\dagger}\rangle$, $\langle\hat{b}_{5}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}}\rangle$ 1
0 $\langle\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{5}}\rangle$ 1 2x1 2x1 2x1 2x1
$\langle\hat{b}_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{6}}\rangle$ 1
$2\gamma_{+}$ $2\alpha_{\mp}$ $\langle\hat{b}_{2}^{2}\rangle$, $\langle\hat{b}_{4}^{2}\rangle$ $\langle\bm{\hat{B}_{1}^{2}}\rangle$ 1 1x4 16x4 1x3 4x3
$\alpha_{-}+\alpha_{+}$ $\langle\hat{b}_{2}\hat{b}_{4}\rangle$ 2 +
$2\alpha_{\mp}$ $\langle\hat{b}_{3}^{2}\rangle$, $\langle\hat{b}_{5}^{2}\rangle$ $\langle\bm{\hat{B}_{2}^{2}}\rangle$ 1 1x4 1x3
$\alpha_{-}+\alpha_{+}$ $\langle\hat{b}_{3}\hat{b}_{5}\rangle$ 2
$2\alpha_{\mp}$ $\langle\hat{b}_{2}^{\dagger 2}\rangle$, $\langle\hat{b}_{4}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{3}^{2}}\rangle$ 1 1x4 1x3
$\alpha_{-}-\alpha_{+}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{4}^{\dagger}\rangle$ 2
$2\alpha_{\mp}$ $\langle\hat{b}_{3}^{\dagger 2}\rangle$, $\langle\hat{b}_{5}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{4}^{2}}\rangle$ 1 1x4 1x3
$\alpha_{-}-\alpha_{+}$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{5}^{\dagger}\rangle$ 2
$2\alpha_{\mp}$ $\langle\hat{b}_{2}\hat{b}_{3}\rangle$, $\langle\hat{b}_{4}\hat{b}_{5}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{2}}\rangle$ 2 2x4 1x4
$\alpha_{+}+\alpha_{-}$ $\langle\hat{b}_{2}\hat{b}_{5}\rangle$, $\langle\hat{b}_{4}\hat{b}_{3}\rangle$ 2
$\alpha_{\mp}+\alpha_{\mp}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{2}\rangle$, $\langle\hat{b}_{4}^{\dagger}\hat{b}_{4}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{1}}\rangle$ 2 2x4 1x4 6x4
$\alpha_{\mp}+\alpha_{\pm}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{4}\rangle$, $\langle\hat{b}_{2}\hat{b}_{4}^{\dagger}\rangle$ 2
$\alpha_{\mp}-\alpha_{\mp}$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{2}\rangle$, $\langle\hat{b}_{5}^{\dagger}\hat{b}_{4}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{1}}\rangle$ 2 2x4 1x4
$\alpha_{\pm}-\alpha_{\mp}$ $\langle\hat{b}_{5}^{\dagger}\hat{b}_{2}\rangle$, $\langle\hat{b}_{3}^{\dagger}\hat{b}_{4}\rangle$ 2
$\alpha_{\mp}-\alpha_{\mp}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{3}\rangle$, $\langle\hat{b}_{4}^{\dagger}\hat{b}_{5}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{2}}\rangle$ 2 2x4 1x4
$\alpha_{\pm}-\alpha_{\mp}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{5}\rangle$, $\langle\hat{b}_{4}^{\dagger}\hat{b}_{3}\rangle$ 2
$\alpha_{\mp}+\alpha_{\mp}$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{3}\rangle$, $\langle\hat{b}_{5}^{\dagger}\hat{b}_{5}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{2}}\rangle$ 2 2x4 1x4
$\alpha_{\mp}+\alpha_{\pm}$ $\langle\hat{b}_{3}^{\dagger}\hat{b}_{5}\rangle$, $\langle\hat{b}_{5}^{\dagger}\hat{b}_{3}\rangle$ 2
$2\alpha_{\mp}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{3}^{\dagger}\rangle$, $\langle\hat{b}_{4}^{\dagger}\hat{b}_{5}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{4}}\rangle$ 2 2x4 1x4
$\alpha_{+}-\alpha_{-}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{5}^{\dagger}\rangle$, $\langle\hat{b}_{4}^{\dagger}\hat{b}_{3}^{\dagger}\rangle$ 2
$\alpha_{\mp}$ $\langle\hat{b}_{2}\hat{b}_{1}\rangle$, $\langle\hat{b}_{4}\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{5}}\rangle$ 2 8x2 16x2 4x2 8x2
$\langle\hat{b}_{2}\hat{b}_{1}^{\dagger}\rangle$ $\langle\hat{b}_{4}\hat{b}_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{1}\hat{B}_{6}}\rangle$ 2
$\langle\hat{b}_{3}\hat{b}_{1}\rangle$, $\langle\hat{b}_{5}\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{5}}\rangle$ 2
$\langle\hat{b}_{3}\hat{b}_{1}^{\dagger}\rangle$ $\langle\hat{b}_{5}\hat{b}_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{2}\hat{B}_{6}}\rangle$ 2
$\alpha_{\mp}$ $\langle\hat{b}_{2}^{\dagger}\hat{b}_{4}\rangle$, $\langle\hat{b}_{4}^{\dagger}\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{5}}\rangle$ 2 8x2 4x2
$\langle\hat{b}_{2}^{\dagger}\hat{b}_{1}^{\dagger}\rangle$ $\langle\hat{b}_{4}^{\dagger}\hat{b}_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{3}\hat{B}_{6}}\rangle$ 2
$\langle\hat{b}_{3}^{\dagger}\hat{b}_{1}\rangle$, $\langle\hat{b}_{5}^{\dagger}\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{5}}\rangle$ 2
$\langle\hat{b}_{3}^{\dagger}\hat{b}_{1}^{\dagger}\rangle$ $\langle\hat{b}_{5}^{\dagger}\hat{b}_{1}^{\dagger}\rangle$ $\langle\bm{\hat{B}_{4}\hat{B}_{6}}\rangle$ 2
0 $\langle\hat{b}_{1}^{2}\rangle$, $\langle\hat{b}_{1}^{\dagger 2}\rangle$ $\langle\bm{\hat{B}_{5}^{2}}\rangle$, $\langle\bm{\hat{B}_{6}^{2}}\rangle$ 1 4x1 4x1 3x1 3x1
$\langle\hat{b}_{1}^{\dagger}\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{6}\hat{B}_{5}}\rangle$ 2
@@ 0 $\langle\hat{b}_{1}\rangle$ $\langle\bm{\hat{B}_{5}}\rangle$ 1 2x1 2x1 2x1 2x1
Table 5: Real and imaginary parts of the complex eigenfrequencies $\bm{\Lambda_{j}^{\rm r}}-i\bm{\Lambda_{j}^{\rm i}}$ of the matrix $\bm{M^{(5)}_{\rm l}}$, given in Eq. (74), for the five-mode linear bosonic system derived from the equations for the FOMs up to second order. We have $\bm{\hat{B}_{1}}\equiv[\hat{b}_{2},\hat{b}_{4}]$, $\bm{\hat{B}_{2}}\equiv[\hat{b}_{3},\hat{b}_{5}]$, $\bm{\hat{B}_{3}}=\bm{\hat{B}_{1}}^{\dagger}$, $\bm{\hat{B}_{4}}=\bm{\hat{B}_{2}}^{\dagger}$, $\bm{\hat{B}_{5}}\equiv[\hat{b}_{1}]$, $\bm{\hat{B}_{6}}=\bm{\hat{B}_{5}}^{\dagger}$, and more details are given in the caption to Tab. 1.